OpenStudy (anonymous):

math help please........8th grade

OpenStudy (anonymous):

fire away

OpenStudy (anonymous):

aWhat is the value of 10C3? A. 30 B. 720 C. 120 D. 1000

OpenStudy (anonymous):

what do you think it is?

OpenStudy (anonymous):

umm b

OpenStudy (anonymous):

incorrect

OpenStudy (anonymous):

10 * 3 * 4 = ?

OpenStudy (anonymous):

120

OpenStudy (anonymous):

Correct! good job!

OpenStudy (welshfella):

a shortcut way of doing these is 10C3 = 10*9*8 ------ 3*2*1 do you see the pattern?

OpenStudy (anonymous):

ohhh yes i do. ok thank ya'll

OpenStudy (anonymous):

Felipe can choose 1 snack from a box containing 3 different granola bars, 4 different kinds of fruit, and 2 different kinds of chips. How many ways can Felipe choose a snack? A. 9 B. 14 C. 5 D. 24 on this one do I just add up all the snacks to get my answer?

OpenStudy (anonymous):

then it would be A?

OpenStudy (welshfella):

no you dont add

OpenStudy (anonymous):

Oh okay can you explain what to do please

OpenStudy (welshfella):

think of a simpler example you have 2 friuts and 3 toffee bars for each of the fruits you can pick any one of 3 toffee bars so how many combinations are there?

OpenStudy (anonymous):

2*3=6?

OpenStudy (welshfella):

fruits are A and B , Bars are C,D and E possible combinations are AC , AD , AE and ?

OpenStudy (welshfella):

right - you multiply

OpenStudy (welshfella):

so extend that to your question..

OpenStudy (anonymous):

so the answer for my problem would be D?

OpenStudy (welshfella):

24 yes

OpenStudy (anonymous):

thank you soo much for explain! :)

OpenStudy (anonymous):

What is the value of (8 – 5)!? A. 40,200 B. 24 C. 5040 D. 6 would this one be D? @welshfella

OpenStudy (welshfella):

yes it = 3! = 3*2*1

OpenStudy (anonymous):

okay yay lol im doing better. i think

OpenStudy (welshfella):

yes keep it up

OpenStudy (anonymous):

How many combinations of 4 students can a teacher choose from 30 students? A. 810,000 B. 27,405 C. 657,720 D. 120

OpenStudy (anonymous):

in this one do i multiply?

OpenStudy (welshfella):

rememnber the shortcut ? 5C3 = 5*3*2 ----- 3*2*2 then 30C4 = 30*29........ -------- 4*3..........

OpenStudy (welshfella):

can you fill in the blanks?

OpenStudy (welshfella):

* correction that is 3*2*1 ( 2 posts ago)

OpenStudy (welshfella):

start from 30 on top of the fraction and multiply by 1 less for a count of 4 then the denominator is 4*3 .... down to 1